OpenStudy (e_s_j_f):

hey, can someone help me with this question? sketch the graph of the function that satisfies the following conditions. a) the domain is [-3,infinity) b) range is (-1,10] c) the curve passes through the point (-3,5) d) the curve is smooth and continious

OpenStudy (e_s_j_f):

i dont understand how to make a curve in this question.

OpenStudy (anonymous):

Domain and range mean values of x and y , respectively and you are given the values to consider in the graph.

OpenStudy (rational):

\[\large f(x) ~~=~~ 11\dfrac{\sqrt{x+3}}{\sqrt{x+k}}-1\] that satisfies \(a,b,d\) conditions we can find the value of \(k\) by using the given point i guess..

OpenStudy (rational):

may be not, as plugging in x=-3 makes that entire fraction 0

OpenStudy (anonymous):

Anyway how did you come up with that equation?

OpenStudy (rational):

domain is restricted to [-3, infty) and the max value has to be 10 so \[f(x) = 10\dfrac{\sqrt{x+3}}{\sqrt{x+3}}\] satisfies those two requirements

OpenStudy (perl):

i can do it more easily with a piecewise defined function

OpenStudy (e_s_j_f):

i understand that, its just the function. i'm not sure about.

OpenStudy (e_s_j_f):

i can do it with a piece wise function as well. how did you come up with the function?

OpenStudy (perl):

im thinking a bell curve (gaussian) function can do the job, in such a manner that it is not defined for x <-3

OpenStudy (perl):

y = 11 exp(-(x+2)^2/2) - 1

OpenStudy (perl):

we can multiply that by sqrt(x+3)^2 / (x+3)

OpenStudy (perl):

$$ \Large { y = \frac{(\sqrt{x+3}~)~^2}{x+3}\cdot 11 e^{-(x+k)^2/2} - 1 } $$ find k such that y(-3) = 5

OpenStudy (rational):

\[y = 5.5\sin(\sqrt{x+3})+4.5\]

OpenStudy (perl):

$$ \LARGE { y = \frac{(\sqrt{x+3}~)~^2}{x+3}\cdot 11 e^{- \frac{ \left( ~x+ 3 - \sqrt{-2\ln\frac{6}{11} } ~ \right) ^2 } {2} } - 1 } $$

OpenStudy (perl):

@rational your range is [-1,10] (i think)

OpenStudy (perl):

the range of sin(sqrt(x+3) is [-1,1] and range 5.5 * [-1,1] + 4.5 = [-1,10] The function I posted suffers from a different drawback, it is not defined at x = -3 but lim x->-3 is 5

OpenStudy (perl):

@E_S_J_F did you find a function that works, piecewise?

OpenStudy (perl):

I think this satisfies all your conditions. $$ \Large { y = \begin{cases} { \frac{(\sqrt{x+3}~)~^2}{x+3}\cdot 11 e^{- \frac{ \left( ~x+ 3 - \sqrt{-2\ln\frac{6}{11} } ~ \right) ^2 } {2} } - 1 ~~\text{if x >-3} \\~\\ ~~~~~~~~~~~~~~~~~~5 ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\text{if x = -3} } \end{cases} } $$

OpenStudy (perl):

THe directions say to graph the function that satisfies the condition, so you don't actually have to find an explicit function.

OpenStudy (rational):

Ahh we need to have an asymptote at bottom

OpenStudy (perl):

right

OpenStudy (perl):

also note that my function is continuous, and differentiable

OpenStudy (perl):

but it would be nice to find a non-piecewise defined function that satisfies all the requirements , explicitly given

OpenStudy (perl):

is there a way to stop a function that is normally defined on R in its domain, but not to exclude that value. The trick i used was to multiply by sqrt(x+3)^2 /(x+3). something similar

OpenStudy (perl):